# On the Sylvester program and Cayley algorithm for vector partition reduction

### v2

@@ Abstract
A vector (compound) partition problem asks for a number of nonnegative integer solutions to a system of several linear Diophantine equations with integer nonnegative coefficients. J.J. Sylvester put forward an idea of reduction of vector partition to a sum of regular (scalar) partitions. For the simplest case of two equations with positive coefficients A. Cayley performed a reduction of the corresponding double partition to a sum of scalar partitions using an algorithm subject to a set of conditions on the coefficients. We suggested [9] a modification of the original Cayley algorithm for the cases when these conditions are not satisfied. In this manuscript we generalize the modified Cayley algorithm to arbitrary number of the Diophantine equations to accomplish the Sylvester program of the vector partition reduction to a combination of scalar partitions.
Keywords: double partition, vector partition.
@@ 1 Introduction
Introducing notation $C[t^{s}](f(t))$ for a coefficient of $t^{s}$ in the expansion of a function $f(t)$ we have
@@
Sylvester proved [12] a statement about splitting of the scalar partition function (SPF) into periodic and non-periodic parts and showed that the partition function may be presented as a sum of ”waves”
@@
Consider a vector partition function $W({\bf s},{\bf D})$ counting the number of integer nonnegative solutions ${\bf x}\geq 0$ to a linear system ${\bf s}={\bf D}\cdot{\bf x}$, where ${\bf D}$ is a nonnegative integer $l\times m$ generator matrix ($l<m$). The vector partition function (VPF) is a natural generalization of scalar partition defined in (2-4) to the vector argument. The generating function for the VPF reads
@@
Several approaches were suggested for VPF computation including the method of residues [1, 13] and geometric decomposition into so called chambers [10] – regions in $l$-dimensional space where the partition function is characterized by a specific expression. The “calculator” computer algebra system was developed to perform derivation of the explicit expressions of vector partitions over chambers [4]. Vector partitions appear to be a powerful tool for the computation of the Kronecker coefficients arising as the structure constants in the decomposition of a tensor product of irreducible representations of the symmetric group into irreducible representations [5]. A method of VPF computation was suggested being a direct generalization of the approach developed in [6] where vector Bernoulli and Eulerian polynomials of higher order were introduced to find explicit expression for $W({\bf s},{\bf D})$ [7]. A drawback of this approach is that it does not provide any mechanism to define the chamber boundaries. As SPF computation requires just known functions (the Bernoulli polynomials of higher order) [6] it is promising to obtain a reduction method expressing a vector partition through a combination of scalar ones.
@@
The first step in this direction was made in the 19th century by J.J. Sylvester [11, 12] who suggested an iterative procedure of reduction of a vector partition into a sum of SPFs; in [11] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of this method applicability [2]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
@@
The author of this manuscript considered the cases when the Cayley formula for double partition fails and showed that its modification successfully solves the problem [9]. Specifically, the elimination of a column with the greatest common divisor (GCD) of its elements $\mbox{gcd}({\bf c}_{i})>1$ (we call it NP-column) leads not to a single SPF but to a weighted sum of such partitions with modified argument. This result is described in Section 3.2 and discussed in Section 3.3. When the second element of a column is zero both the original Cayley method and its generalization cannot be applied so that this case requires special treatment. In Section 3.4 we prove that the contribution of such a column to the double partition is zero.
@@
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to each column. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. Three examples of triple partition reduction are given in Appendices.
@@
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 where we show that in this case a vector partition reduces to a convolution of scalar partitions. Other details of computational implementation of the reduction step are presented in Sections 6.1 and 6.2 for P-column and NP-column respectively. We show in Section 6.3 that the reduction of a triple partition always produces NP-columns. This implies that the reduction of any VPF for a matrix with $l>2$ rows in most of the cases leads to application of the extended version of the Cayley algorithm. Additional details of the iterative reduction procedure are discussed in Section 6.4 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
The problem of vector (compound) partitions has a long history and J.J. Sylvester made a significant contribution to its solution. Sylvester in 1858 wrote: ”Any given system of simultaneous simple equations to be solved in positive integers being proposed, the determination of the number of solutions of which they admit may in all cases be made to depend upon the like determination for one or more systems of equations of a certain fixed standard form. When a system of $r$ equations between $n$ variables of the aforesaid standard form is given, the determination of the number of solutions in positive integers of which it admits may be made to depend on the like determination for
@@
In other words, Sylvester claimed that $r$-tuple vector partition for $n$ variables can be reduced to a sum of $\binom{n}{r-1}$ scalar partitions and the reduction algorithm is an iterative process based on the variable elimination. Sylvester considered a specific double partition problem as an illustration of his method and determined regions (chambers) on a plane each having a unique expression for VPF valid in this region only. He showed that the expressions in the adjacent chambers coincide at their common boundary (see also [10]).
@@ 3 Cayley method for double partitions
Computation of vector partition $W({\bf s},{\bf D})$ in (7) is performed by iterative elimination of $(l-1)$ variables $t_{k},(2\leq k\leq l)$ [1]. Each elimination step includes partial fraction expansion (PFE) w.r.t. the eliminated variable with subsequent coefficient evaluation. This step is equivalent to elimination of $k$-th row of augmented $l\times(m+1)$ matrix ${\bf E}=\{{\bf s}={\bf c}_{0},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of matrix ${\bf D}$ and vector ${\bf s}$; the same time the numerical columns of ${\bf D}$ are eliminated one by one. The number of new $(l-1)\times m$ matrices is equal to $m$. This algorithm was employed in [2] for a two-row positive matrix (see also [11]). The original Cayley algorithm and its generalization are exhaustively discussed in [9] and below we present their main points.
@@ 3.1 Original Cayley algorithm
$A_{1}(x,x^{-b_{1}/\beta_{1}})=\prod_{i\neq 1}^{m}(1-x^{b_{i}-b_{1}\beta_{i}/\beta_{1}})^{-1}.$ (9)
It can be written as
@@
where $A_{1k}$ are rational functions in $x$.
Consider evaluation of the contribution $C_{r,\rho}^{1}$ of the term $T_{1}(x,y)$, it is equal to the coefficient
@@
$C_{r,\rho}^{1}$ $=$ $C[x^{r}y^{\rho}]\left(\frac{A_{1}(x,y)}{1-x^{b_{1}/\beta_{1}}y}\right)=C[x^{r}y^{\rho}]\left(\frac{A_{1}(x,x^{-b_{1}/\beta_{1}})}{1-x^{b_{1}/\beta_{1}}y}\right)$
$=$ $C[x^{r}]\left(x^{\rho b_{1}/\beta_{1}}A_{1}(x,x^{-b_{1}/\beta_{1}})\right)=C[x^{r-\rho b_{1}/\beta_{1}}]\left(A_{1}(x,x^{-b_{1}/\beta_{1}})\right)$
$=$ $C[x^{r\beta_{1}-\rho b_{1}}]\left(A_{1}(x^{\beta_{1}},x^{-b_{1}})\right).$
@@
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W^{2}_{i}({\bf s})=\sum_{i=1}^{m}W(L_{i},{\bf d}_{i}),\quad L_{i}=r\beta_{i}-b_{i}\rho,\quad d_{ij}=b_{j}\beta_{i}-b_{i}\beta_{j},\ j\neq i.$ (13)
@@
we observe that $L_{i}$ and elements $d_{ij}$ of ${\bf d}_{i}$ are given by determinants of ${\bf D}_{i0}$ and ${\bf D}_{ij}$, respectively. Noncollinearity of the columns ${\bf c}_{i}$ implies that all elements in the generator sets ${\bf d}_{i}$ are nonzero, but some of them might be negative (say, $d_{ij_{k}}<0$ for $1\leq k\leq K_{i}$). Noting that $(1-t^{-a})^{-1}=-t^{a}(1-t^{a})^{-1}$ we find an equivalent scalar partition with positive generators only
$W(L_{i},{\bf d}_{i})=(-1)^{K}W(L_{i}+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|),\quad|{\bf d}_{i}|=\{|d_{ij}|\}.$ (15)
The partition $W(L_{i}+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|)$ in r.h.s. of (15) corresponds to the Diophantine equation
$\sum_{j\neq i}^{m}x_{j}|d_{ij}|=L_{i}({\bf s})=s_{1}\beta_{i}-s_{2}b_{i}+\sum_{k=1}^{K_{i}}d_{ij_{k}},$
where $K_{i}$ integers $d_{ij_{k}}<0$ are negative and $x_{j}\geq 0$ are nonnegative integers. When $L_{i}({\bf s})$ is negative this equation has no solutions and we can rewrite (15) as
$W^{2}_{i}({\bf s})=W(L_{i},{\bf d}_{i})=A_{i}W(L_{i}({\bf s}),|{\bf d}_{i}|),\quad A_{i}=(-1)^{K_{i}}H(L_{i}({\bf s})),$ (16)
@@
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W(L^{\prime}_{i},{\bf d}^{\prime}_{i}),\quad L^{\prime}_{i}=-L_{i},\quad{\bf d}^{\prime}=-{\bf d}.$ (17)
It should be noted that $L^{\prime}_{i}({\bf s})\neq-L_{i}({\bf s})$ and the individual contributions $W(L_{i},{\bf d}_{i})$ and $W(L^{\prime}_{i},{\bf d}^{\prime}_{i})$ differ. This means that the order of the rows of ${\bf E}$ does not influence the value of the vector partition $W({\bf s},{\bf D})$ but it does affect individual summands $W^{2}_{i}({\bf s})$ in (16).
@@ 3.2 Cayley method generalization
When Cayley method fails we suggested an alternative approach [9] based on explicit computation of $A_{1}(x,y)$. Turning to $S_{1}(x,y_{0})$ write it as
@@
with $10\leq k_{i}\leq\beta_{1}-$ 1. Expanding r.h.s. of (18) we obtain
@@
$x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}}y_{0}^{\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}\rfloor}y_{0}^{(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})\bmod{\beta_{1}}}$
@@
$\bar{W}_{i}^{2}({\bf s})=\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}A_{i,{\bf j}}W_{i}^{2}(({\bf s}-{\bf j})/\beta_{i}),\ \ A_{i,{\bf j}}=(-1)^{K_{i}}a_{i,{\bf j}}H(L_{i}(({\bf s}-{\bf j})/\beta_{i}))\delta_{j_{2},s_{2}\bmod{\beta_{i}}}.$ (26)
3.3 General Cayley method for P- and NP-columns

### v3

@@ Abstract
A vector partition problem asks for a number of nonnegative integer solutions to a system of linear Diophantine equations with integer nonnegative coefficients. J.J. Sylvester suggested to use variable elimination for the reduction of vector partition to a sum of scalar partitions. In the simplest case of two equations with positive coefficients A. Cayley performed a reduction of the corresponding double partition to a sum of scalar partitions using the algorithm subject to a set of conditions on the coefficients. We suggested a modification of the original Cayley algorithm for the cases when these conditions are not satisfied [14]. In this manuscript we generalize the modified Cayley algorithm to arbitrary number of the Diophantine equations to accomplish the Sylvester program of the vector partition reduction to a combination of scalar partitions.
Keywords: integer partition, double partition, vector partition.
@@ 1 Introduction
Introducing a notation $C[t^{s}](f(t))$ for a coefficient of $t^{s}$ in the expansion of a function $f(t)$ we have
@@
Sylvester proved [17] a statement about splitting of the scalar partition into periodic and non-periodic parts and showed that the partition function may be presented as a sum of ”waves”
@@
Consider a vector partition function $W({\bf s},{\bf D})$ counting the number of integer nonnegative solutions ${\bf x}\geq 0$ to a linear system ${\bf s}={\bf D}\cdot{\bf x}$, where ${\bf D}$ is a nonnegative integer $l\times m$ generator matrix ($l<m$). The vector partition is a natural generalization of scalar partition defined in (2-4) to the vector argument. The generating function for the vector partition reads
@@
An alternative method of vector partition computation was suggested being a direct generalization of the approach developed in [11] where vector Bernoulli and Eulerian polynomials of higher order were introduced to find explicit expression for $W({\bf s},{\bf D})$ [12]. A drawback of this approach is that it does not provide any mechanism to define the chamber boundaries. As scalar partition computation requires just known functions (the Bernoulli polynomials of higher order) [11] it is promising to obtain a reduction method expressing a vector partition through a combination of scalar ones.
@@
The first step in this direction was made in the 19th century by J.J. Sylvester [16, 17] who suggested an iterative procedure of reduction of a vector partition into a sum of scalar partitions; in [16] he gave an example of such a reduction for a double partition. Based on this approach A. Cayley developed an algorithm of double partition reduction and established a set of conditions of this method applicability [4]. Cayley showed that each column ${\bf c}_{i}$ of the two-row matrix ${\bf D}$ gives rise to a scalar partition and elements of this column must be relatively prime (we call it P-column). Another limitation of the Cayley method is that the columns ${\bf c}_{i}$ should represent noncollinear vectors. An application of the Cayley algorithm to the computation of the Gaussian polynomial coefficients is considered in [13]. A short historical excurse into the (seemingly forgotten) ideas of Sylvester and Cayley is given in Section 2 while the essence of the original Cayley algorithm is presented in Section 3.1.
@@
The author of this manuscript considered the case when the Cayley formula for double partition fails and discussed its modification that successfully resolves the problem [14]. Specifically, the elimination of a column with the greatest common divisor (GCD) of its elements $\mbox{gcd}({\bf c}_{i})>1$ (we call it NP-column) leads not to a single scalar partition but to a weighted sum of such partitions with modified argument. This result is described in Section 3.2 and discussed in Section 3.3.
@@
In Section 4 we extend the reduction algorithm to $l$-tuple (for arbitrary $l\geq 2$) vector partition defined by a $l\times m$ nonnegative generator matrix ${\bf D}$ made of $m$ columns ${\bf c}_{i}$. The first step of the algorithm discussed in Section 4.1 is the partial fraction expansion of the generating function (6) into $m$ terms corresponding to elimination of each of $m$ unknowns. The contribution of a column is computed in Section 4.2 and it is either a single partition for P-columns (Section 4.2.1) or a superposition of such partitions (Section 4.2.2) for NP-column. When the last element of a column is zero both the original Cayley method and its generalization cannot be applied directly so that this case requires a special treatment. In Section 4.3 we prove that the contribution of such a column to the vector partition is zero.
@@
The case of collinear columns of the generator matrix ${\bf D}$ is addressed in Section 5 where we show that the vector partition reduces to a convolution of scalar partitions. Additional details of the iterative reduction procedure are discussed in Section 6 while the concluding remarks are presented in Section 7.
@@ 2 Sylvester-Cayley method
The problem of vector (compound) partitions has a long history that started in 1768 when Euler announced the generating function (6) for double partition ($l=2$) [7] and followed by J.J. Sylvester presenting an alternative approach in 1858. Sylvester wrote: ”Any given system of simultaneous simple equations to be solved in positive integers being proposed, the determination of the number of solutions of which they admit may in all cases be made to depend upon the like determination for one or more systems of equations of a certain fixed standard form. When a system of $r$ equations between $n$ variables of the aforesaid standard form is given, the determination of the number of solutions in positive integers of which it admits may be made to depend on the like determination for
@@
In other words, Sylvester claimed that $r$-tuple vector partition for $n$ variables can be reduced to a sum of $\binom{n}{r-1}$ scalar partitions and the reduction algorithm is an iterative process based on the variable elimination. Sylvester considered a specific double partition problem as an illustration of his method and determined regions (chambers) on a plane each having a unique expression for vector partition valid in this region only. He showed that the expressions in the adjacent chambers coincide at their common boundary (see also [15]).
@@ 3 Cayley method for double partitions
Computation of vector partition $W({\bf s},{\bf D})$ in (7) is performed by iterative elimination of $(l-1)$ variables $t_{k},(2\leq k\leq l)$ [2]. Each elimination step includes partial fraction expansion (PFE) w.r.t. the eliminated variable with subsequent coefficient evaluation. This step is equivalent to elimination of $k$-th row of augmented $l\times(m+1)$ matrix ${\bf E}=\{{\bf s}={\bf c}_{0},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of the matrix ${\bf D}$ and vector ${\bf s}$; the same time the numerical columns of ${\bf D}$ are eliminated one by one. The number of new $(l-1)\times m$ matrices is equal to $m$. This algorithm was employed in [4] for a two-row positive matrix (see also [16]). The original Cayley algorithm and its generalization are exhaustively discussed in [14] and below we present their main points.
@@ 3.1 Original Cayley algorithm
$A_{1}(x,y_{0})=\prod_{i\neq 1}^{m}(1-x^{b_{i}}y_{0}^{\beta_{i}})^{-1}=\prod_{i\neq 1}^{m}(1-x^{b_{i}-b_{1}\beta_{i}/\beta_{1}})^{-1}.$ (9)
Use the relation $1/(1-x)=1/(1-x^{n})\sum_{k=0}^{n-1}x^{k}$ to write $A_{1}(x,y_{0})$ as
@@
where $A_{1k}$ are rational functions in $x$. Consider evaluation of the contribution $C_{r,\rho}^{1}$ of the term $T_{1}(x,y)$, it is equal to the coefficient
@@
$C_{r,\rho}^{1}$ $=$ $C[x^{r}y^{\rho}]\left(\frac{A_{1}(x,x^{-b_{1}/\beta_{1}})}{1-x^{b_{1}}y^{\beta_{1}}}\right)=C[x^{r}]\left(x^{\rho b_{1}/\beta_{1}}A_{1}(x,x^{-b_{1}/\beta_{1}})\right)$ (11)
$=$ $C[x^{r-\rho b_{1}/\beta_{1}}]\left(A_{1}(x,x^{-b_{1}/\beta_{1}})\right)=C[x^{r\beta_{1}-\rho b_{1}}]\left(A_{1}(x^{\beta_{1}},x^{-b_{1}})\right).$
@@
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W^{2}_{i}({\bf s})=\sum_{i=1}^{m}W(L_{i}({\bf s}),{\bf d}_{i}),\quad L_{i}({\bf s})=r\beta_{i}-b_{i}\rho,\quad d_{ij}=b_{j}\beta_{i}-b_{i}\beta_{j},\ j\neq i.$ (13)
@@
we observe that $L_{i}({\bf s})$ and elements $d_{ij}$ of ${\bf d}_{i}$ are given by determinants of ${\bf D}_{i0}$ and ${\bf D}_{ij}$, respectively. Noncollinearity of the columns ${\bf c}_{i}$ implies that all elements in the generator sets ${\bf d}_{i}$ are nonzero, but some of them might be negative (say, $d_{ij_{k}}<0$ for $1\leq k\leq K_{i}$). Noting that $(1-t^{-a})^{-1}=-t^{a}(1-t^{a})^{-1}$ we find an equivalent scalar partition with positive generators only
$W(L_{i}({\bf s}),{\bf d}_{i})=(-1)^{K}W(L_{i}({\bf s})+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|),\quad|{\bf d}_{i}|=\{|d_{ij}|\}.$ (15)
The partition $W(L_{i}({\bf s})+\sum_{k=1}^{K_{i}}d_{ij_{k}},|{\bf d}_{i}|)$ in the r.h.s. of (15) corresponds to the Diophantine equation
$\sum_{j\neq i}^{m}x_{j}|d_{ij}|=L^{\prime}_{i}({\bf s})=s_{1}\beta_{i}-s_{2}b_{i}+\sum_{k=1}^{K_{i}}d_{ij_{k}},$
where $K_{i}$ integers $d_{ij_{k}}<0$ are negative and $x_{j}\geq 0$ are nonnegative integers. When $L^{\prime}_{i}({\bf s})$ is negative this equation has no solutions and we can rewrite (15) as
$W^{2}_{i}({\bf s})=W(L_{i}({\bf s}),{\bf d}_{i})=A_{i}W(L^{\prime}_{i}({\bf s}),|{\bf d}_{i}|),\quad A_{i}=(-1)^{K_{i}}H(L^{\prime}_{i}({\bf s})),$ (16)
@@
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W(\tilde{L}_{i}({\bf s}),\tilde{\bf d}_{i}),\quad\tilde{L}_{i}({\bf s})=-L_{i}({\bf s}),\quad\tilde{\bf d}=-{\bf d}.$ (17)
It should be noted that $\tilde{L}^{\prime}_{i}({\bf s})\neq-L^{\prime}_{i}({\bf s})$ and the individual contributions $W(L_{i}({\bf s}),{\bf d}_{i})$ and $W(\tilde{L}_{i}({\bf s}),\tilde{\bf d}_{i})$ differ. This means that the order of the rows of ${\bf E}$ does not influence the value of the vector partition $W({\bf s},{\bf D})$ but it does affect individual summands $W^{2}_{i}({\bf s})$ in (16).
It is worth to repeat after Dickson [6] that the Cayley algorithm corresponds to the elimination of each variable $x_{i}$ from the system ${\bf s}={\bf D}\cdot{\bf x}$ producing $m$ equations $L_{i}={\bf d}_{i}\cdot\bm{x}^{\prime}_{i}$ where $\bm{x}^{\prime}_{i}=\{x_{1},x_{2},\ldots,x_{i-1},x_{i+1},,\ldots,x_{m}\}$ is obtained by dropping $x_{i}$ from ${\bf x}$.
@@ 3.2 Cayley method generalization
When the Cayley method fails we suggested an alternative approach [14] based on explicit computation of $A_{1}(x,y)$. Turning to $S_{1}(x,y_{0})$ in (10) write it as
@@
with $10\leq k_{i}\leq\beta_{1}-$ 1. Expanding the r.h.s. of (18) we obtain
@@
$x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}}y_{0}^{\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}}=x^{\bm{K}^{\prime}_{p}\cdot\bm{b}^{\prime}_{1}-b_{1}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})/\beta_{1}\rfloor}y_{0}^{(\bm{K}^{\prime}_{p}\cdot\bm{\beta}^{\prime}_{1})\bmod{\beta_{1}}}\;,$
@@
$\bar{W}_{i}^{2}({\bf s})=\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}A_{i,{\bf j}}W_{i}^{2}(({\bf s}-{\bf j})/\beta_{i}),\ \ A_{i,{\bf j}}=(-1)^{K_{i}}a_{i,{\bf j}}H(L^{\prime}_{i}(({\bf s}-{\bf j})/\beta_{i}))\delta_{j_{2},s_{2}\bmod{\beta_{i}}}.$ (26)
3.3 Cayley method for P- and NP-columns